You are planning to buy a cellphone with a good quality in order to attend the online class. The average price of 50 cellphones is Php13,600 with a margin of error of Php 283.75 and a confidence level of 90%. Show solution if necessary






What is the lower limit?





What is the upper limit?





What is the confidence interval?





If you have F13.200.00, is the amount lies within the confidence interval?





The length of confidence interval is?

1
Expert's answer
2022-04-24T17:45:10-0400

lower limit =sample mean -margin error


lower limit =13600-283.75


lower limit =13316.25




Upper Limit=sample mean +margin error


Upper Limit =13600+283.75


Upper Limit =13883.75




90% Confidence interval =(Lower Limit ,Upper Limit )


90% Confidence interval =(13316.25,13883.75)

Length of interval is 13883.75-13316.25=567.5

The amount P 13200 Iies outside the confidence interval